Find the magnitude of the dot product of the two vectors B=6i+j+2k,D=-i+3j.what is the angle between B and D.
1
Expert's answer
2020-01-02T10:22:26-0500

"B=6i+j+2k \newline D=-i+3j \newline
B.D=(6i+j+2k ).(-i+3j) \newline
B.D=(-6+3)=-3
\\~\\"

From the equation,

"B.D=|B||D|cos(\theta)" where "\theta" is the angle between two vectors

"\cos(\theta)=\frac{-3}{(\sqrt{41})(\sqrt{10})} \newline
\theta=cos^{-1}(\frac{-3}{\sqrt{410}}) \newline
\theta=98.52\degree"